OpenStudy (anonymous):

Which pair of ratios forms a proportion? 10/15 & 5/30 8/40 & 16/20 18/13 & 9/26 18/12 & 36/24 I think it is either 10/15 & 5/30 or 18/12 & 36/24

hero (hero):

Set the ratios equal to each other, then cross multiply

OpenStudy (anonymous):

18/12 & 36/24

OpenStudy (anonymous):

18/12 and 36/24 are in proportion.

OpenStudy (anonymous):

Thank you guys :)

OpenStudy (diyadiya):

10/15=2/3 5/30=1/6 so they are not in proportion

OpenStudy (anonymous):

Cheers!!!!!!!!!!!!!!!!!